# GPU Acceleration of SQL Analytics on Compressed Data

### v1

@@ Abstract.
Work done during an internship at the Microsoft Gray Systems Lab while affiliated with Columbia University.
@@ 1. Introduction
The rest of the paper is organized as follows. Section 2 briefly describes the tensor approach used by TQP that we leverage in this work. Section 3 presents the encoding schemes for compression that we use. Section 4 introduces our novel and parallel implementations of core transformations on compressed data. Sections 58 show how we use them in novel ways to efficiently implement relational operators on GPUs. Section 9 presents experimental results and speedups with our techniques. Section 10 discusses related work and Section 11 concludes the paper.
2. Background: TQP
The Tensor Query Processor (TQP) (He et al., 2022; Asada et al., 2022) converts SQL queries into tensor programs, then runs them using Tensor Computation Runtimes (TCRs), such as PyTorch, on hardware backends such as GPUs (He et al., 2022; Gandhi et al., 2022; Cao et al., 2023) and APUs (Cui et al., 2023). This makes TQP easily portable to different devices while at the same time benefiting from the optimizations provided by the TCRs. In this work, we continue to use TQP’s approach to get the same portability and performance advantages, but extend it to show how we can convert queries to tensor programs that can operate on lightweight-compressed data.
@@ 3.1. Basic Encodings
  • Plain: This is the existing tensor representation used by TQP as we discussed in Section 2. For every column, there is a 1:1 mapping from a row position in the column to the corresponding position in the tensor representing the values in the column.
  • RLE: We use three tensors to represent a set of RLE runs: (1) values ($v$), (2) start positions ($s$), and (3) end positions ($e$). Tensors $v$, $s$, and $e$ are of equal length, and the $i^{th}$ entry represents a run with value $v_{i}$ spanning row numbers in the closed and continuous interval [$s_{i}$, $e_{i}$]. The positions are zero-based and unique. Tensors $v$, $s$, $e$ are sorted by $s$ (equivalently, by $e$). Note that, we could have used a tensor for run lengths ($l$) instead of for end positions ($e$). We can choose either lengths or end positions based on convenience, and calculate the other at run time using the equation: $l=e-s+1$. Runs are non-overlapping by position.
  • Index: We use two tensors to represent index encoding: (1) values ($v$), and (2) positions ($p$). Tensors $v$ and $p$ are of equal length, sorted by $p$, and the $i^{th}$ entry represents value $v_{i}$ at row number $p_{i}$. The positions are zero-based and unique.
While Plain encoding has an implicit 1:1 correspondence between row numbers and positions in the tensor representation, RLE and Index encoding explicitly track row numbers. This also allows for efficient representation when there are gaps in the underlying data, e.g., when some portions of the column are de-selected after application of filter predicates. Thus, for RLE, for two consecutive entries at tensor positions $i$ and $i+1$, we have $s_{i+1}\geq e_{i}+1$. Similarly, for Index encoding, we have $p_{i+1}\geq p_{i}+1$. In contrast, Plain encoding does not allow for gaps in the tensor representation.
@@ 3.2. Composite Encodings
  • Plain + Index: PyTorch requires a uniform data type for all elements within a single tensor. However, outliers can force the entire tensor to use a larger data type than necessary. To enable bit-width reduction with narrower data types, we can combine Plain with Index encoding. We store most values in one PyTorch tensor with a narrower type and represent the outliers separately using Index encoding.
@@
Plain $\rightarrow$ Plain + Index example: When handling a column with values $[1,2,3,10^{10},10^{10}]$, the outliers at indices 4 and 5 would force the entire tensor to use a larger data type. We separate the data: regular values in Plain tensor $[1,2,3]$ and outliers in a separate value tensor $[10^{10},10^{10}]$ with their positions in index tensor $[4,5]$. In a dataset with 1 billion elements, if a few outliers require int64 but most values fit in int8, Plain representation would force all values into int64, using 8 GiB of memory (8 bytes × 1 billion). With Plain + Index, we store most values in int8 (1 GiB) plus a small overhead for outliers and their indices, saving $\sim$7 GiB of memory while preserving all data values.
@@ 3.3. Data and Mask column representations
  • The domain of values for masks is restricted to {$T,F$}.
@@
  • In position-explicit encodings (RLE, Index) for masks, the value tensor ($v$) is not needed, since all tracked values are implicitly equal to $T$.
@@ 4. Primitives
Due to space limitation, in this section we include pseudo-code only for a few primitives, notably the intersection primitives that are heavily used for implementing relational operators. Functions in bold font are provided by PyTorch and have been optimized by PyTorch developers for different backends including GPUs. Note that there are no explicit for loops or conditional checks in the codes, which helps to maximize utilization of the available hardware parallelism.
@@ 4.1. Range Intersection
Our approach is inspired by bioinformatics techniques for chromosome range intersections (Layer et al., 2013), as detailed in Algorithm 1. The ”range_intersect” algorithm efficiently computes the intersection of two RLE inputs, $c_{1}$ and $c_{2}$, by first bucketizing the start ($c_{1}.s$) and end ($c_{1}.e$) points of the first input relative to the second (Lines 1-2). It then counts intersections (Line 3) and generates index tensors $idx_{1}$ and $idx_{2}$ for each intersecting range (Lines 4-6), using the helper function range_arange (Algorithm 2) for $idx_{2}$. Unlike PyTorch’s arange, which operates on single start/length values, range_arange accepts tensors of starts and lengths to generate multiple concatenated sequences. Finally, the algorithm determines the start and end points of the intersections (Lines 7), producing the resulting tensors $s$ and $e$. To optimize performance, the input with the fewer ranges should always be used as $c_{1}$.
@@
1:$bin_{s}$leftarrow bucketize($c_{1}.s$, $c_{2}.e$, right=False)
2:$bin_{e}$leftarrow bucketize($c_{1}.e$, $c_{2}.s$, right=True)
3:cnt $\leftarrow$ $bin_{e}$ $bin_{s}$
@@
5:$idx_{1}$leftarrow repeat_interleave(arange, cnt)
6:$idx_{2}$leftarrow range_arange($bin_{s}$, cnt)
7:$s,e$leftarrow max($c_{1}.s$[$idx_{1}$], $c_{2}.s$[$idx_{2}$]), min($c_{1}.e$[$idx_{1}$], $c_{2}.e$[$idx_{2}$])
@@
Consider the range intersection between the positional ranges of two RLE masks, $c_{1}$ and $c_{2}$, as shown in Figure 2. Mask $c_{1}$ has a single range represented as $c_{1}.s=[2]$, $c_{1}.e=[7]$. Mask $c_{2}$ has three ranges represented as $c_{2}.s=[1,4,6]$, $c_{2}.e=[3,5,8]$. Recall that masks only represent ranges corresponding to a value of $T$.
To compute their intersection using the range_intersect algorithm (Algorithm 1):
  • Step 1: Bucketize start positions. We bucketize the start position ($c_{1}.s$) of Mask $c_{1}$ relative to the end positions ($c_{2}.e$) of Mask $c_{2}$. $bin\_s=\textbf{bucketize}([2],[3,5,8])=[0]$. This means the start point 2 comes before the first end point 3.
  • Step 2: Bucketize end positions. We bucketize the end positions ($c_{1}.e$) of Mask $c_{1}$ relative to the start positions ($c_{2}.s$) of Mask $c_{2}$. $bin\_e=\textbf{bucketize}([7],[1,4,6],right=True)=[3]$. This means the end point $7$ comes after all start positions in $c_{2}$.
  • Step 3: Count overlaps. We count the overlaps by subtracting $bin\_s$ from $bin\_e$. $cnt=bin\_e-bin\_s=[3]-[0]=[3]$. This indicates that the range in $c_{1}$ overlaps with all 3 ranges in $c_{2}$.
[…]
  • Step 5: Compute intersection points. We compute the actual intersection points by taking the maximum of start points and minimum of end points. $s=\textbf{max}(c_{1}.s[idx\_1],c_{2}.s[idx\_2])=\textbf{max}([2,2,2],[1,4,6])=[2,4,6]$ and $e=\textbf{min}(c_{1}.e[idx\_1],\\ c_{2}.e[idx\_2])=\textbf{min}([7,7,7],[3,5,8])=[3,5,7]$.
So, the final output is an RLE mask with $s=[2,4,6]$ and $e=[3,5,7]$, representing exactly the segments where both input masks $c_{1}$ and $c_{2}$ have True values.
@@ 4.2. Index Intersection
These operations handles intersections for Index encoded data ($c.p$). A common task is finding which positions ($c_{idx}.p$) from an Index list fall within any RLE ranges ($c_{rle}.s,c_{rle}.e$). We provide two algorithms for this: idx_in_rle (Algorithm 3) and rle_contain_idx (Algorithm 5). The main computational work involves the bucketize operation performed by each. idx_in_rle uses bucketize($c_{idx}.p$, $c_{rle}.s$), while rle_contain_idx uses bucketize($c_{rle}.s$, $c_{idx}.p$) and bucketize($c_{rle}.e$, $c_{idx}.p$). The choice between them for optimal performance depends on the relative input sizes: idx_in_rle is generally preferred when $|c_{idx}.p|\ll|c_{rle}.s|$. We also provide idx_in_idx (Algorithm 4) for intersecting two Index lists.
@@
Let us consider idx_in_rle with an Index column $c_{idx}$ and an RLE column $c_{rle}$. Suppose $c_{idx}.p=[2,4,7]$ and $c_{rle}.s=[0,6]$, $c_{rle}.e=[2,7]$. We want to find which elements in $c_{idx}.p$ fall within the RLE ranges [0-2] or [6-7] defined by $c_{rle}$.
  • Step 1: Bucketize positions. bucketize($c_{idx}.p$, $c_{rle}.s$, right=True) gives $[1,1,2]$. Subtracting 1 yields $bin=[0,0,1]$. This tells us which RLE range each position $p$ might belong to (position 2 maps to range 0, 4 maps to range 0, 7 maps to range 1).
  • Step 2: Verify containment. We check two conditions for each position $p$ in $c_{idx}.p$: (1) bin $\geq$ 0 ensures the position isn’t before the first start $s$ in $c_{rle}.s$, and (2) $p\leq c_{rle}.e$[bin] ensures the position is not past the end $e$ of its assigned RLE range. For position 2: bin=0. $0\geq 0$ ($T$) and $22\leq c_{rle}.e[0]=$ 2 ($T$). 2 is included. For position 4: bin=0. $0\geq 0$ ($T$) and $24\leq c_{rle}.e[0]=$ 2 ($F$). 4 is excluded. For position 7: bin=1. $1\geq 0$ ($T$) and $77\leq c_{rle}.e[1]=$ 7 ($T$). 7 is included.
  • Step 3: Apply mask. The resulting mask is $[T,F,T]$. Applying this to $c_{idx}.p$ gives the final result $p_{out}$ = [2, 7].
@@
1:bin $\leftarrow$ bucketize($c_{idx}.p$, $c_{rle}.s$, right=True)
@@
3:mask $\leftarrow$ (bin $\geq$ 0) $\land$ ($c_{idx}.p$leq $c_{rle}.e$[bin])
4:return $c_{idx}.p$[mask]
@@
1:bin $\leftarrow$ bucketize($c_{1}.p$, $c_{2}.p$, right=True) - 1
2:mask $\leftarrow$ (bin $\geq$ 0) $\land$ ($c_{1}.p=c_{2}.p[bin]$)
3:return $c_{1}.p$[mask]
Example 4 (rle_contain_idx).
Let us illustrate rle_contain_idx using the same input data as the idx_in_rle example for direct comparison. Suppose $c_{idx}.p=[2,4,7]$ and the RLE column $c_{rle}$ has ranges defined by $c_{rle}.s=[0,6]$ and $c_{rle}.e=[2,7]$. The ranges are [0-2] and [6-7]. We want to find which positions in $c_{idx}.p$ are contained within any of these RLE ranges.
  • Step 1: Bucketize RLE starts. We bucketize each RLE start position ($c_{rle}.s$) against the sorted index positions ($c_{idx}.p$). $bin_{s}=\textbf{bucketize}(c_{rle}.s,c_{idx}.p)=\textbf{bucketize}([0,6],[2,4,7])=[0,2]$. (0 is before index 0 value ’2’; 6 is before index 2 value ’7’).
  • Step 2: Bucketize RLE ends. We bucketize each RLE end position ($c_{rle}.e$) against $c_{idx}.p$, using $right=True$ and subtracting 1. $bin_{e}=\textbf{bucketize}(c_{rle}.e,c_{idx}.p$, right=True) - 1 = bucketize([2, 7], [2, 4, 7], right=True) - 1 = [1, 3] - 1 = [0, 2]. (2 is $\leq$ index 0 value 2, giving index 1; 7 is $\leq$ index 2 value 7, giving index 3. Then subtract 1).
  • Step 4: Apply mask. We apply $mask$ to $bin_{s}$ and $bin_{e}$. They remain unchanged: $bin_{s}=[0,2]$, $bin_{e}=[0,2]$.
  • Step 5: Convert RLE indices to position indices. We use rle_to_index to generate the flat list of indices into $c_{idx}.p$ that correspond to the ranges defined by the masked $bin_{s}$ and $bin_{e}$. For range 1 ($s=0,e=0$): indices [0]. For range 2 ($s=2,e=2$): indices [2]. Combined indices: $[0,2]$. Let this result be $idx_{flat}$.
  • Step 6: Gather positions. We select the positions from $c_{idx}.p$ using $idx_{flat}$: $p_{out}=c_{idx}.p[idx_{flat}]=c_{idx}.p[[0,2]]=[2,7]$.
The final output is $p_{out}=[2,7]$, which are exactly the positions from $c_{idx}.p$ that fall within the RLE ranges [0-2] or [6-7]. This matches the output of the idx_in_rle example for the same input.
Finally, the idx_in_idx algorithm (Algorithm 4) works similarly to idx_in_rle, using bucketize to find potential matches between two sorted position lists ($c_{1}.p,c_{2}.p$) and then verifying exact equality.
@@
1:$bin_{s}$leftarrow bucketize($c_{rle}.s$, $c_{idx}.p$)
2:$bin_{e}$leftarrow bucketize($c_{rle}.e$, $c_{idx}.p$, right=True) - 1
@@
4:$bin_{s},bin_{e}\leftarrow$ $bin_{s}$[mask], $bin_{e}$[mask]
5:return $c_{idx}.p$[rle_to_index($bin_{s}$, $bin_{e}$)]

### v2

@@ 1. Introduction
The rest of the paper is organized as follows. Section 2 briefly describes TQP and the PyTorch primitives that we leverage in this work. Section 3 presents the encoding schemes for compression that we use. Section 4 introduces our novel and parallel implementations of core transformations on compressed data. Sections 58 show how we use them in novel ways to efficiently implement relational operators on GPUs. Section 9 presents experimental results and speedups with our techniques. Section 10 discusses related work and Section 11 concludes the paper.
2. Background
We first describe TQP, our query execution framework, then introduce the PyTorch primitives used in our compressed operators.
2.1. TQP
The Tensor Query Processor (TQP) (He et al., 2022; Asada et al., 2022) converts SQL queries into tensor programs, then runs them using Tensor Computation Runtimes (TCRs), such as PyTorch (Paszke et al., 2019), on hardware backends such as GPUs (He et al., 2022; Gandhi et al., 2022; Cao et al., 2023) and APUs (Cui et al., 2023). This makes TQP easily portable to different devices while at the same time benefiting from the optimizations provided by the TCRs. In this work, we continue to use TQP’s approach to get the same portability and performance advantages, but extend it to show how we can convert queries to tensor programs that can operate on lightweight-compressed data.
@@ 2.1. TQP
2.2. PyTorch Primitives
Tensors are the fundamental data structures in PyTorch. While tensors can be high-dimensional (for ML applications), in this work we use them as one-dimensional arrays that directly correspond to database columns. Our algorithms rely on several PyTorch primitives that have been highly optimized for ML tasks but proves effective for database query processing.
Tensor Indexing []: The bracket operator supports three modes: (1) single indexing (e.g., $data[0]$ selects the element at position 0), (2) multi-indexing with index tensors (e.g., $data[[0,2]]$ selects elements at positions 0 and 2), and (3) filtering with boolean masks (e.g., $data[mask]$ returns elements where mask is True). For example, given $data=[10,20,30,40]$: $data[[0,2]]$ returns $[10,30]$; $data[[True,False,True,False]]$ also returns $[10,30]$.
[…]
arange(end): Generates a tensor with values from 0 to end1end-1. For example, $arange(5)$ returns $[0,1,2,3,4]$.
repeat_interleave(input, repeats): Repeats each element of inputinput consecutively by the corresponding count in repeatsrepeats. For example, repeat_interleave([10,20],[2,1])repeat\_interleave([10,20],[2,1]) returns $[10,10,20]$.
unique(input): Returns unique elements and inverse indices mapping original elements to their positions. For example, $unique([3,1,3,2])$ returns unique $[1,2,3]$ and inverse indices $[2,0,2,1]$.
scatter(src, index, reduce): Accumulates values from srcsrc at positions specified by indexindex using the reducereduce operation. Multiple values targeting the same index are combined according to the reducereduce. For example, $scatter([10,20,30],[0,1,0],reduce=sum)$ results in $[40,20]$ (10+30 at index 0, 20 at index 1).
@@ 3.1. Basic Encodings
We will discuss multiple columns, each with multiple fields for their tensor representations. For notation clarity: we use subscripts to distinguish different columns (e.g., $c_{1}$, $c_{2}$), and superscripts for field access within each column’s encoding (e.g., $c_{1}^{start}$, $c_{2}^{pos}$).
@@
  • Plain: This is the existing tensor representation used by TQP as we discussed in Section 2.1. For every column, there is a 1:1 mapping from a row position in the column to the corresponding position in the tensor representing the values in the column.
  • RLE: We use three tensors to represent a set of RLE runs: (1) values ($c^{val}$), (2) start positions ($c^{start}$), and (3) end positions ($c^{end}$). Tensors $c^{val}$, $c^{start}$, and $c^{end}$ are of equal length, and the $i^{th}$ entry represents a run with value $c^{val}[i]$ spanning row numbers $c^{start}[i]\rightarrow c^{end}[i]$. The positions are zero-based and unique. Tensors $c^{val}$, $c^{start}$, $c^{end}$ are sorted by $c^{start}$ (equivalently, by $c^{end}$). Note that, we could have used a tensor for run lengths ($c^{len}$) instead of for end positions ($c^{end}$). We can choose either lengths or end positions based on convenience, and calculate the other at run time using the equation: $c^{len}=c^{end}-c^{start}+1$. Runs are non-overlapping by position.
  • Index: We use two tensors to represent index encoding: (1) values ($c^{val}$), and (2) positions ($c^{pos}$). Tensors $c^{val}$ and $c^{pos}$ are of equal length, sorted by $c^{pos}$, and the $i^{th}$ entry represents value $c^{val}[i]$ at row number $c^{pos}[i]$. The positions are zero-based and unique.
While Plain encoding has an implicit 1:1 correspondence between row numbers and positions in the tensor representation, RLE and Index encoding explicitly track row numbers. This also allows for efficient representation when there are gaps in the underlying data, e.g., when some portions of the column are deselected after application of filter predicates. Thus, for RLE, for two consecutive entries at tensor positions $i$ and $i+1$, we have $c^{start}_{i+1}\geq c^{end}_{i}+1$. Similarly, for Index encoding, we have $c^{pos}_{i+1}\geq c^{pos}_{i}+1$. In contrast, Plain encoding does not allow for gaps in the tensor representation.
@@ 3.2. Composite Encodings
  • Plain + Index: PyTorch requires a uniform data type for all elements within a single tensor. However, outliers can force the entire tensor to use a larger data type than necessary. To enable bit-width reduction with narrower data types, we can combine Plain with Index encoding. We store most values in one PyTorch tensor with a narrower type and represent the outliers separately using Index encoding. At positions corresponding to outliers, the Plain tensor contains uninitialized PyTorch tensor values (values happened to be in that memory location). These uninitialized values are never used during query execution.
@@
Plain $\rightarrow$ Plain + Index example: When handling a column with values $1,2,3,10^{10},10^{10}]$, the outliers at indices 3 and 4 would force the entire tensor to use a larger data type. We separate the data: regular values in Plain tensor $[1,2,3]$ and outliers in a separate value tensor $10^{10},10^{10}]$ with their positions in index tensor $[3,4]$. In a dataset with 1 billion elements, if a few outliers require int64 but most values fit in int8, Plain representation would force all values into int64, using 8 GiB of memory (8 bytes × 1 billion). With Plain + Index, we store most values in int8 (1 GiB) plus a small overhead for outliers and their indices, saving $\sim$7 GiB of memory while preserving all data values.
@@ 3.3. Data and Mask column representations
  • The domain of values for masks is restricted to {T,FT,F}.
@@
  • In position-explicit encodings for masks, the value tensor ($v$) is not needed, since all tracked values are implicitly equal to $T$.
@@ 4. Primitives
Due to space limitation, in this section we include pseudo-code only for a few primitives, notably the intersection primitives that are heavily used for implementing relational operators. Functions in bold font are provided by PyTorch and have been optimized by PyTorch developers for different backends including GPUs. Note that there are no explicit for loops or conditional checks in the codes, which helps to maximize utilization of the GPU parallelism.
@@ 4.1. Range Intersection
Our approach is inspired by bioinformatics techniques for chromosome range intersections (Layer et al., 2013), as detailed in Algorithm 1. The ”range_intersect” algorithm efficiently computes the intersection of two RLE inputs, $c_{1}$ and $c_{2}$, by first bucketizing the start ($c_{1}^{start}$) and end ($c_{1}^{end}$) points of the first input relative to the second (Lines 1-2). It then counts intersections (Line 3) and generates index tensors $idx_{1}$ and $idx_{2}$ for each intersecting range (Lines 4-6), using the helper function range_arange (Algorithm 2) for $idx_{2}$. Unlike PyTorch’s arange, which operates on single start/length values, range_arange accepts tensors of starts and lengths to generate multiple concatenated sequences. Finally, the algorithm determines the start and end points of the intersections (Line 7), producing the resulting tensors $s$ and $e$. To optimize performance, the input with the fewer ranges should always be used as $c_{1}$.
@@
[…]
3:cnt $\leftarrow$ $bin_{e}$ - $bin_{s}$
@@
[…]
7:$s,e$ $\leftarrow$ max($c_{1}^{start}$$idx_{1}$], $c_{2}^{start}$$idx_{2}$]), min($c_{1}^{end}$$idx_{1}$], $c_{2}^{end}$$idx_{2}$])
@@
Consider the range intersection between the positional ranges of two RLE masks, $c_{1}$ and $c_{2}$, as shown in Figure 2. Mask $c_{1}$ has a single range represented as $c_{1}^{start}=[2]$, $c_{1}^{end}=[7]$. Mask $c_{2}$ has three ranges represented as $c_{2}^{start}=[1,4,6]$, $c_{2}^{end}=[3,5,8]$. Recall that masks only represent ranges corresponding to a value of $T$.
To compute their intersection using range_intersect (Algorithm 1):
  • Step 1: Bucketize start positions. We bucketize the start position ($c_{1}^{start}$) of Mask $c_{1}$ relative to the end positions ($c_{2}^{end}$) of Mask $c_{2}$. $bin_{s}=\textbf{bucketize}([2],[3,5,8])=[0]$. This means the start point 2 comes before the first end point 3.
  • Step 2: Bucketize end positions. We bucketize the end positions ($c_{1}^{end}$) of Mask $c_{1}$ relative to the start positions ($c_{2}^{start}$) of Mask $c_{2}$. $bin_{e}=\textbf{bucketize}([7],[1,4,6],right=True)=[3]$. This means the end point $7$ comes after all start positions in $c_{2}$.
  • Step 3: Count overlaps. We count the overlaps by subtracting $bin_{s}$ from $bin_{e}$. $cnt=bin_{e}-bin_{s}=[3]-[0]=[3]$. This indicates that the range in $c_{1}$ overlaps with all 3 ranges in $c_{2}$.
  • Step 4: Create index tensors. We create index tensors that identify which ranges from each mask participate in each overlap. $idx_{1}=\textbf{repeat\_interleave}([0],[3])=[0,0,0]$ (indices into $c_{1}$) and $idx_{2}=\textbf{range\_arange}([0],[3])=[0,1,2]$ (indices into $c_{2}$).
  • Step 5: Compute intersection points. We compute the actual intersection points by taking the maximum of start points and minimum of end points. $s=\textbf{max}(c_{1}^{start}[idx_{1}],c_{2}^{start}[idx_{2}])=\textbf{max}([2,2,2],[1,4,6])=[2,4,6]$ and $e=\textbf{min}(c_{1}^{end}[idx_{1}],\\ c_{2}^{end}[idx_{2}])=\textbf{min}([7,7,7],[3,5,8])=[3,5,7]$.
The output is an RLE mask with $s=[2,4,6]$ and $e=[3,5,7]$, representing the segments where both $c_{1}$ and $c_{2}$ have True values.
@@ 4.2. Index Intersection
These operations handle intersections for Index encoded data ($c^{pos}$). A common task is finding which positions ($c_{idx}^{pos}$) from an Index list fall within any RLE ranges ($c_{rle}^{start},c_{rle}^{end}$). We provide two algorithms for this: idx_in_rle (Algorithm 3) and rle_contain_idx (Algorithm 5). The main computational work involves the bucketize operation performed by each. idx_in_rle uses bucketize($c_{idx}^{pos}$, $c_{rle}^{start}$), while rle_contain_idx uses bucketize($c_{rle}^{start}$, $c_{idx}^{pos}$) and bucketize($c_{rle}^{end}$, $c_{idx}^{pos}$). The choice between them for optimal performance depends on the relative input sizes: idx_in_rle is generally preferred when $|c_{idx}^{pos}|\ll|c_{rle}^{start}|$. We also provide idx_in_idx (Algorithm 4) for intersecting two Index lists.
@@
Let us consider idx_in_rle with an Index column $c_{idx}$ and an RLE column $c_{rle}$. Suppose $c_{idx}^{pos}=[2,4,7]$ and $c_{rle}^{start}=[0,6]$, $c_{rle}^{end}=[2,7]$. We want to find which elements in $c_{idx}^{pos}$ fall within the RLE ranges [0-2] or [6-7] defined by $c_{rle}$.
  • Step 1: Bucketize positions. bucketize($c_{idx}^{pos}$, $c_{rle}^{start}$, right=True) gives $[1,1,2]$. Subtracting 1 yields $bin=[0,0,1]$. This tells us which RLE range each position $p$ might belong to (position 2 maps to range 0, 4 maps to range 0, 7 maps to range 1).
  • Step 2: Verify containment. We check two conditions for each position $p$ in $c_{idx}^{pos}$: (1) bin $\geq$ 0 ensures the position is not before the first start $s$ in $c_{rle}^{start}$, and (2) $p\leq c_{rle}^{end}$[bin] ensures the position is not past the end $e$ of its assigned RLE range. For position 2: bin=0. $0\geq 0$ ($T$) and $2\leq c_{rle}^{end}[0]=2$ ($T$). 2 is included. For position 4: bin=0. $0\geq 0$ ($T$) and $4\leq c_{rle}^{end}[0]=2$ ($F$). 4 is excluded. For position 7: bin=1. $1\geq 0$ ($T$) and $7\leq c_{rle}^{end}[1]=7$ ($T$). 7 is included.
  • Step 3: Apply mask. The resulting mask is $[T,F,T]$. Applying this to $c_{idx}^{pos}$ gives the final result $p_{out}$ = [2, 7].
@@
1:bin $\leftarrow$ bucketize($c_{idx}^{pos}$, $c_{rle}^{start}$, right=True)
@@
3:mask $\leftarrow$ (bin $\geq$ 0) $\land$ ($c_{idx}^{pos}$ $\leq$ $c_{rle}^{end}$[bin])
4:return $c_{idx}^{pos}$[mask]
@@
1:bin $\leftarrow$ bucketize($c_{1}^{pos}$, $c_{2}^{pos}$, right=True) - 1
2:mask $\leftarrow$ (bin $\geq$ 0) $\land$ ($c_{1}^{pos}=c_{2}^{pos}[bin]$)
3:return $c_{1}^{pos}$[mask]
@@
[…]
@@
[…]
5:return $c_{idx}^{pos}$[rle_to_index($bin_{s}$, $bin_{e}$)]
Example 4 (rle_contain_idx).
Let us illustrate rle_contain_idx using the same input data as the idx_in_rle example for direct comparison. Suppose $c_{idx}^{pos}=[2,4,7]$ and the RLE column $c_{rle}$ has ranges defined by $c_{rle}^{start}=[0,6]$ and $c_{rle}^{end}=[2,7]$. The ranges are [0-2] and [6-7]. We want to find which positions in $c_{idx}^{pos}$ are contained within any of these RLE ranges.
  • Step 1: Bucketize RLE starts. We bucketize each RLE start position ($c_{rle}^{start}$) against the sorted index positions ($c_{idx}^{pos}$). $bin_{s}=\textbf{bucketize}(c_{rle}^{start},c_{idx}^{pos})=\textbf{bucketize}([0,6],[2,4,7])=[0,2]$. (0 is before index 0 value ’2’; 6 is before index 2 value ’7’).
  • Step 2: Bucketize RLE ends. We bucketize each RLE end position ($c_{rle}^{end}$) against $c_{idx}^{pos}$, using $right=True$ and subtracting 1. $bin_{e}=\textbf{bucketize}(c_{rle}^{end},c_{idx}^{pos}$, right=True) - 1 = bucketize([2, 7], [2, 4, 7], right=True) - 1 = [1, 3] - 1 = [0, 2]. (2 is $\leq$ index 0 value 2, giving index 1; 7 is $\leq$ index 2 value 7, giving index 3. Then subtract 1).
  • Step 4: Apply mask. We apply maskmask to $bin_{s}$ and $bin_{e}$. They remain unchanged: $bin_{s}=[0,2]$, $bin_{e}=[0,2]$.
  • Step 5: Convert RLE indices to position indices. We use rle_to_index to generate the flat list of indices into $c_{idx}^{pos}$ that correspond to the ranges defined by the masked $bin_{s}$ and $bin_{e}$. For range 1 ($s=0,e=0$): indices [0]. For range 2 ($s=2,e=2$): indices [2]. Combined indices: $[0,2]$. Let this result be $idx_{flat}$.
  • Step 6: Gather positions. We select the positions from $c_{idx}^{pos}$ using $idx_{flat}$: $p_{out}=c_{idx}^{pos}[idx_{flat}]=c_{idx}^{pos}[[0,2]]=[2,7]$.
The final output is $p_{out}=[2,7]$, which are exactly the positions from $c_{idx}^{pos}$ that fall within the RLE ranges [0-2] or [6-7]. This matches the output of the idx_in_rle example for the same input.
Finally, the idx_in_idx algorithm (Algorithm 4) similarly to idx_in_rle uses bucketize to find potential matches between two sorted position lists ($c_{1}^{pos},c_{2}^{pos}$) and then verifying exact equality.
4.3. Primitive Microbenchmarks
[…]
(a) range_intersect
(b) idx_in_rle
(c) idx_in_idx
(d) rle_contain_idx
[…]